y=6x+8
y=-6x-16
algebra 1
substitution

Answer by stanbon(75887)

y=6x+8
y=-6x-16
algebra 1
substitution
-----------------
6x+8 = -6x-16
12x = -24
x = -2
-----------
y = 6*-2+8 = -12+8 = -4
===========================
Cheers,
Stan H.